OpenStudy (anonymous):

the mass of jupiter is about 1.9x10^24 metric tons. The mass of the sun is about 1.998x10^27 metric tons. About how many times the mass of Jupiter is the mass of the sun? Give answer in scientific notation.

OpenStudy (anonymous):

Approx. 10^(3) i dont know scientific notations sorry

OpenStudy (stormfire1):

\[\frac{1.998 * 10^{27}}{1.9*10^{24}}=1.051*10^3=1.1 * 10^3\](corrected for significant digits)